OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is perpendicular to the line 2x -3y = 12 and passes through the point (2, 6).

OpenStudy (mathlegend):

First put the equation in the form of y=mx+b

OpenStudy (mathlegend):

-3y=-2x+12 -- ------- -3 -3 y = 2/3x - 4

OpenStudy (mathlegend):

Remember Perpendicular ... is reciprocal that is negative

OpenStudy (mathlegend):

y-6=-3/2(x-2) y-6=-3/2x+3 +6 +6 --------------- y = -3/2x + 9

OpenStudy (mathlegend):

@raerae01

OpenStudy (anonymous):

Thank you! So makes sense now! (: